OpenStudy (anonymous):

how do I state the slope of 3x-4y=12

OpenStudy (anonymous):

First solve it in y=mx+b form. -4y=-3x+12. y=3/4x-3. The slope is the term attached to x, called m in y=mx+b. aka -3!

hero (hero):

Add 4y to both sides: 3x = 12 + 4y Subtract 12 from both sides: 3x - 12 = 4y Divide both sides by 4: (3/4)x - 3 = y Since we have the form y = mx + b, m = 3/4, thus 3/4 is the slope.

OpenStudy (anonymous):

I meant to write 3/4 ^^ sorry wasn't paying attention to that bit

OpenStudy (anonymous):

@_mia_rachel that's fine thank you both for your help